QUESTION 22 If the Marginal Product of capital is 6 and the Marginal Product of labor is 3; the prices of capital and labor are $10 and $2 respectively. What should the manager do? Increase output Substitute in more labor for less capital Pay workers less Substitute in more capital for less labor
Answer #1

Answer

the firm should hire the input at the marginal product per dollar is equal

marginal product per dollar =marginal input /price of input

marginal product per dollar of labor =3/2=1.5 per dollar

marginal product per dollar of capital =6/10=0.6 per dollar

the MP per dollar is higher for labor so the firm should increase labor and decrease capital up to the MU per dollar of both is equal.